Write the standard form of the equation of the parabola with the following:
1. Center (0,0) and r=7
2. Center (-2, 4) and r=4

Answer by lwsshak3(11628)

Write the standard form of the equation of the parabola with the following:
1. Center (0,0) and r=7
2. Center (-2, 4) and r=4
**
I think you mean circles instead of parabolas.
Standard form of equation for circles: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
..
1. Center (0,0) and r=7
x^2+y^2=7^2
x^2+y^2=49
..
2. Center (-2, 4) and r=4
(x+2)^2+(y-4)^2=4^2
(x+2)^2+(y-4)^2=16